# Strong-lensing cosmography using third-generation gravitational-wave detectors

### v1

@@ 4 Effect of lens distribution
Figure 2: Posteriors (90% credible regions) of cosmological parameters that illustrates the bias in the inference due to the use of wrong models of the HMF. In the legends, BB, BJ, BS and BT denote different scenarios when our simulated observation of lensed signals is produced using Behroozi model, and then cosmological parameters are recovered using Behroozi, Jenkins, SMT and Tinker08 models, respectively. Left panel: Results corresponding to a merger rate $5\times 10^{5}\;\mathrm{yr}^{-1}$. We can see that true cosmology is not recovered within 95% credible region except for the case of BB. However, the precision for all cases remains almost the same. Right panel: Results corresponding to a merger rate $5\times 10^{4}\;\mathrm{yr}^{-1}$. Here, we can see that the amount of bias is relatively small due to the decreased precision of the cosmological parameters.
@@
The distribution of the redshift and other parameters of the lenses are determined using the HMF models (see section 2.2). Here we investigate how selecting the incorrect model for the lens distribution (both redshift and other parameters) can affect the inference of cosmological parameters. We broadly follow the same steps as outlined in section 3. However, here we assume redshift distribution given by Dominik [48] to be the one presented by nature, and we will vary the lens distribution while keeping the other parameters fixed. We consider four different models of HMF, namely Behroozi [53], Jenkins [55], SMT [56], and Tinker08 [57]. We simulate $N$ lensed events with time delays $\{\Delta t_{i}\}_{i=1}^{N}$ using the Behroozi model. $N$ is drawn from a Poisson distribution with mean $\Lambda(\vec{\Omega}_{\mathrm{true}},T_{\mathrm{obs}})$, where $\vec{\Omega}_{\mathrm{true}}$ is the assumed “true” value of cosmological parameters. This is our simulated observational data. We then estimate the cosmological parameters using all of these HMF models.
@@
We observe that there is a bias in the inference of cosmological parameters when we choose the wrong model of lens distribution. Figure 2 shows an example set of posteriors (90% credible regions) on the cosmological parameters when different HMF models are used in the parameter estimation. The left (right) panel shows results corresponding to an assumed merger rate $R=5\times 10^{5}\;\mathrm{yr}^{-1}$ ($5\times 10^{4}\;\mathrm{yr}^{-1}$) for an observation time of $10\;\mathrm{yrs}$. In the left panel the true cosmology is not recovered within 95% credible region except when the true HMF is used in the parameter estimation, while amount of bias is relatively small in the right panel due to the decreased precision.
@@
The posteriors in figure 2 show one realisation of the observation scenario. Each of the posterior could be randomly shifted due to the finite number of observed events (Poisson fluctuations). To get the statistical nature of the biases, we perform this analysis using $\sim 10^{3}$ catalogs of observations of same statistical nature. The so-called probability-probability (p-p) plots that shows in what fraction of the simulated observations, the true values are recovered within a given credible interval (figure 3). If the theoretical models agree with the data, the p-p plot should show a diagonal line. We see that diagonal p-p plots are obtained only when the simulated HMF model is used in the parameter estimation.
@@
Figure 3: The cumulative distribution of the quantiles (p-p plots) in which the true cosmological parameters are recovered from $\sim 10^{3}$ recovery tests. The legend box shows the scenarios considered (same as figure 2). Any deviation from the diagonal line suggest there is a bias in the recovery. The the larger the deviation, the greater is the bias in the inference. Solid (dashed) lines show the recovery for $\Omega_{m}$ ($H_{0}$). In the left panel, the results corresponds to a merger rate of $5\times 10^{5}\;\mathrm{yr}^{-1}$ show that there are biases, except in the BB sceinario. In the right panel, the results for a merger rate of $5\times 10^{4}\;\mathrm{yr}^{-1}$ indicate that there is relatively lower bias due to the decreased precision in the estimation of the cosmological parameters.
@@
While this is a cause of concern, we show below that Bayesian model selection generally allows us to identify the correct model of the HMF. We compute the ratios of Bayesian evidences [equation (2)] of different HMF models and show that the right HMF model is almost always preferred. We simulate lensed events with time delays using model B (Behroozi) and then recover the cosmological parameters using all these models. We observe that for the model that consistently shows bias in its parameter estimates, the evidence is smaller than the “true” model. The ratio of the evidences (the Bayes factor) between the true and false models is greater than one for most of the simulations. For instance, in the parameter estimation using the J (Jenkins) model typically cause large bias in the estimated parameters (see figure 2) . However, the Bayes factor between B and J models is higher than one over $\sim 80\%$ of the time, indicating that the true model is generally preferred (figure 4). In the parameter estimation using the T (Tinker) model, the true model is preferred only for $\sim 50\%$ of the time; however, the bias in the estimated parameters using the T model is generally smaller.
@@
Figure 4: The cumulative distribution of $\log_{10}$ (Bayes factor) between the “right” (B) and “wrong” (J, S or T) HMF models, computed from $\sim 10^{3}$ simulated catalogs. The results corresponds to merger rates of $5\times 10^{5}\;\mathrm{yr}^{-1}$ (left) and $5\times 10^{4}\;\mathrm{yr}^{-1}$ (right). We can see that HMF models that cause significant bias in the parameter estimation have low Bayes factors most of the time.
@@ 5 Effect of contamination
We assume that the total number $N_{\mathrm{tot}}$ of detected BBH mergers observed follow a Poisson distribution of mean $\Lambda_{\mathrm{tot}}$ while the detected number of strongly lensed mergers $N$ follows a Poisson distribution of mean $\Lambda$. We define the contamination fraction $\kappa$ as the ratio between the expected number of falsely identified lensed pairs and the number of truly identified lensed pairs
@@
where $u\equiv\Lambda/\Lambda_{\mathrm{tot}}$ is the expected lensing fraction. Note that $\kappa$ is a function of cosmological parameters and the observing period $T_{\mathrm{obs}}$ as the lensing fraction depends on $\vec{\Omega}$ and the total number of observed events depends on $T_{\mathrm{obs}}$. For simplicity of notation, we don’t explicitly write down its dependence on $\vec{\Omega}$ and $T_{\mathrm{obs}}$. The contamination fraction also depends on $k_{0}\equiv\alpha/\epsilon$, which depends on the receiver operating characteristic (ROC) of a given lensing identification method. This can be estimated by performing the same analysis on simulated lensed and unlensed GW events (see, e.g. [40]). To keep the contamination fraction low (about $10\%$), we would need $\alpha\sim 10^{-9}$ for $\Lambda\sim 10^{6}$ and $u=0.01$. This is likely to be achievable in the future GW observations due to the increased precision of measurements.
@@
and use them for the cosmological parameter estimation.
To gauge the effect of contamination on the cosmological parameter inference, we simulate lensed events following the Dominick redshift distribution, assuming the true cosmology $\Omega_{\mathrm{true}}=\{\Omega_{m}=0.316,H_{0}=67.3\}$, with $R=5\times 10^{5}$ and $T_{\mathrm{obs}}=10\mathrm{yrs}$. This corresponds to a true expected detection of $\Lambda_{\mathrm{tot}}=5\times 10^{6}$ binaries. We assume that the lensing identification method has a false positive probability of $\alpha=10^{-9}$. We assume different true positive probabilities: $\epsilon=0.4,0.5,0.6,0.8$, corresponding to $k_{0}\equiv\alpha/\epsilon=[2.5,2,1.67,0.8]\times 10^{-9}$. Figure 5 shows the posteriors on the cosmological parameters from an analysis that takes into the effect of contamination. When there is no contamination $k_{0}=0$, we reproduce the results from figure 1. Additionally, data contamination only worsens the precision of our measurement, without causing any systematic biases (see p-p plot in figure 5). Table 2 tabulates the expected precision in the measurement of cosmological parameters with different levels of contamination.
No contamination $\epsilon=0.4$ $\epsilon=0.5$ $\epsilon=0.6$ $\epsilon=0.8$
@@
Figure 5: Left panel: Posteriors on the cosmological parameters (90% credible regions) when we assume different amounts of data contamination. The legends show $k_{0}\equiv\alpha/\epsilon$ in units of $10^{-9}$, where $\alpha$ ($\epsilon$) is the false (true) positive probability of the lensing identification algorithm. Data contamination only worsens the precision of the posteriors, without causing any systematic biases. Right panel: p-p plot obtained from $\sim 10^{3}$ recovery tests for $\alpha=10^{-9}$ and $\epsilon=0.5$. This indicates no systematic bias in cosmological parameter recovery when we account for contamination.
@@ 6 Conclusions and future work
In this paper we presented a detailed exposition of a statistical method for cosmography from the observation of a large number ($\sim 10^{4}$) of strongly lensed BBH mergers observable by XG detectors. This method, first presented in [1], compares the observed number of strongly lensed GW events and their time delay distribution (between lensed images) with observed events to infer cosmological parameters. We showed that the prospective constraints from the XG detectors are comparable to the current best measurements, but probing a cosmological epoch ($z\sim 1-10$) that is not probed by other observations. This can potentially shed some light on the “tensions” that exist between current cosmological observations [65], some hinting the failure of the flat $\Lambda$CDM model.
The observed number of lensed events and their time delay distributions will depend on the distribution of GW sources as well as lenses, apart from the cosmological parameters. Properties of the GW source distribution could be measured accurately from the large number ($\sim 10^{6}$) of unlensed GW signals that will dominate the data. Since the distribution of the GW source properties is currently largely unknown, we considered a few different astrophysical models for the same, and showed that the expected constraints don’t vary significantly. We also showed that the limited ability of our data analysis algorithms to distinguish between lensed and unlensed GW events, resulting in some amount contamination in the sample of lensed GW events, will not bias our inference. This can be avoided by incorporating the effect of data contamination in our Bayesian likelihood models.
@@
While we focus on BBH mergers in this paper, observations of strongly lensed binaries involving neutron stars could also be powerful. Since BNS could be observed only out to a smaller redshifts ($z\lesssim 2$ even for XG detectors), their lensing optical depth will be smaller compared to BBHs which could be observed out to very large redshifts ($z\sim 10-100$). However, BNS mergers are expected to be large in number ($\sim 10^{6}$ detections per year in XG [66]) and, owing to their long signal duration in the detector band, their properties could be measured with much better accuracy. Furthermore, some them could also produce observable EM counterparts. Strongly lensed BNSs could be very powerful in probing low-redshift cosmology. We are exploring this in an ongoing work. In conclusion, future observations of strongly lensed GWs will enable several powerful new probes of cosmology.
@@ Acknowledgments:
We thank Otto Akseli Hannuksela for his careful review of the manuscript and useful comments. We are also grateful to the members of the ICTS Astrophysical Relativity group and the LIGO-Virgo-KAGRA collaboration’s working group on gravitational lensing for useful discussions. Our research is supported by the Department of Atomic Energy, Government of India, under Project No. RTI4001. Numerical calculations were performed using the Alice cluster at ICTS.

### v2

@@ 5 Effect of lens distribution
Figure 3: Posteriors (95% credible regions) of cosmological parameters that illustrate the bias in the inference due to the use of wrong models of the HMF. In the legends, BB, BJ, BS and BT denote different scenarios when our simulated observation of lensed signals is produced using Behroozi model, and then cosmological parameters are recovered using Behroozi, Jenkins, SMT and Tinker08 models, respectively. Left panel: Results corresponding to a merger rate $5\times 10^{5}\;\mathrm{yr}^{-1}$. We can see that true cosmology is not recovered within 95% credible region except for the case of BB. However, the precision for all cases remains almost the same. Right panel: Results corresponding to a merger rate $5\times 10^{4}\;\mathrm{yr}^{-1}$. Here, we can see that the amount of bias is relatively small due to the decreased precision of the cosmological parameters.
@@
The distribution of the redshift and other parameters of the lenses are determined using the HMF models (see section 2.2). Here we investigate how selecting the incorrect model for the lens distribution (both redshift and other parameters) can affect the inference of cosmological parameters. We broadly follow the same steps as outlined in section 3. However, here we assume the redshift distribution given by Dominik [47] to be the one presented by nature, and we will vary the lens distribution while keeping the other parameters fixed. We consider four different models of HMF, namely Behroozi [51], Jenkins [54], SMT [55], and Tinker08 [56]. We simulate $N$ lensed events with time delays $\{\Delta t_{i}\}_{i=1}^{N}$ using the Behroozi model. $N$ is drawn from a Poisson distribution with mean $\Lambda(\vec{\Omega}_{\mathrm{true}},T_{\mathrm{obs}})$, where $\vec{\Omega}_{\mathrm{true}}$ is the assumed “true” value of cosmological parameters. This is our simulated observational data. We then estimate the cosmological parameters using all of these HMF models.
@@
We observe that there is a bias in the inference of cosmological parameters when we choose the wrong model of lens distribution. Figure 3 shows an example set of posteriors (95% credible regions) on the cosmological parameters when different HMF models are used in the parameter estimation. The left (right) panel shows results corresponding to an assumed merger rate $R=5\times 10^{5}\;\mathrm{yr}^{-1}$ ($5\times 10^{4}\;\mathrm{yr}^{-1}$) for an observation time of $10\;\mathrm{yrs}$. In the left panel the true cosmology is not recovered within 95% credible region except when the true HMF is used in the parameter estimation, while the amount of bias is relatively small in the right panel due to the decreased precision.
@@
The posteriors in figure 3 show one realisation of the observation scenario. Each of the posterior could be randomly shifted due to the finite number of observed events (Poisson fluctuations). To get the statistical nature of the biases, we perform this analysis using $\sim 10^{3}$ catalogs of observations of same statistical nature. The so-called probability-probability (p-p) plots show in what fraction of the simulated observations, the true values are recovered within a given credible interval (figure 4). If the theoretical models agree with the data, the p-p plot should show a diagonal line. We see that diagonal p-p plots are obtained only when the simulated HMF model is used in the parameter estimation.
@@
Figure 4: The cumulative distribution of the quantiles (p-p plots) in which the true cosmological parameters are recovered from $\sim 10^{3}$ recovery tests. The legend box shows the scenarios considered (same as figure 3). Any deviation from the diagonal line suggest there is a bias in the recovery. The larger the deviation, the greater is the bias in the inference. Solid (dashed) lines show the recovery for $\Omega_{m}$ ($H_{0}$). In the left panel, the results corresponding to a merger rate of $5\times 10^{5}\;\mathrm{yr}^{-1}$ show that there are biases, except in the BB scenario. In the right panel, the results for a merger rate of $5\times 10^{4}\;\mathrm{yr}^{-1}$ indicate that there is relatively lower bias due to the decreased precision in the estimation of the cosmological parameters.
@@
While this is a cause of concern, we show below that Bayesian model selection generally allows us to identify the correct model of the HMF. We compute the ratios of Bayesian evidences [equation (2)] of different HMF models and show that the right HMF model is almost always preferred. We simulate lensed events with time delays using model B (Behroozi) and then recover the cosmological parameters using all these models. We observe that for the model that consistently shows bias in its parameter estimates, the evidence is smaller than the “true” model. The ratio of the evidences (the Bayes factor) between the true and false models is greater than one for most of the simulations. For instance, in the parameter estimation using the J (Jenkins) model typically causes large bias in the estimated parameters (see figure 3) . However, the Bayes factor between B and J models is higher than one over $\sim 80\%$ of the time, indicating that the true model is generally preferred (figure 5). In the parameter estimation using the T (Tinker) model, the true model is preferred only for $\sim 50\%$ of the time; however, the bias in the estimated parameters using the T model is generally smaller.
@@
Figure 5: The cumulative distribution of $\log_{10}$ (Bayes factor) between the “right” (B) and “wrong” (J, S or T) HMF models, computed from $\sim 10^{3}$ simulated catalogs. The results correspond to merger rates of $5\times 10^{5}\;\mathrm{yr}^{-1}$ (left) and $5\times 10^{4}\;\mathrm{yr}^{-1}$ (right). We can see that HMF models that cause significant bias in the parameter estimation have low Bayes factors most of the time.
@@ 6 Effect of contamination
We assume that the total number $N_{\mathrm{tot}}$ of detected BBH mergers observed follows a Poisson distribution of mean $\Lambda_{\mathrm{tot}}$ while the detected number of strongly lensed mergers $N$ follows a Poisson distribution of mean $\Lambda$. We define the contamination fraction $\kappa$ as the ratio between the expected number of falsely identified lensed pairs and the number of truly identified lensed pairs
@@
where $u\equiv\Lambda/\Lambda_{\mathrm{tot}}$ is the expected lensing fraction. Note that $\kappa$ is a function of cosmological parameters and the observing period $T_{\mathrm{obs}}$ as the lensing fraction depends on $\vec{\Omega}$ and the total number of observed events depends on $T_{\mathrm{obs}}$. For simplicity of notation, we don’t explicitly write down its dependence on $\vec{\Omega}$ and $T_{\mathrm{obs}}$. The contamination fraction also depends on $k_{0}\equiv\alpha/\epsilon$, which depends on the receiver operating characteristic (ROC) of a given lensing identification method. This can be estimated by performing the same analysis on simulated lensed and unlensed GW events (see, e.g. [40]). To keep the contamination fraction low (about $10\%$), we would need $\alpha\sim 10^{-9}$ for $\Lambda\sim 10^{6}$ and $u=0.01$. This is likely to be achievable in future GW observations due to the increased precision of measurements.
@@
and use them for the cosmological parameter estimation. Above, $p_{\mathrm{lens}}(\Delta t~{}|~{}\vec{\Omega},T_{\mathrm{obs}})$ is given by equation (6), while
$p_{\mathrm{unlens}}(\Delta t~{}|~{}T_{\mathrm{obs}})\propto(T_{\mathrm{obs}}-\Delta t)\Theta(T_{\mathrm{obs}}-\Delta t).$ (24)
To gauge the effect of contamination on the cosmological parameter inference, we simulate lensed events following the Dominick redshift distribution, assuming the true cosmology $\Omega_{\mathrm{true}}=\{\Omega_{m}=0.316,H_{0}=67.3\}$, with $R=5\times 10^{5}$ and $T_{\mathrm{obs}}=10\;\mathrm{yrs}$. This corresponds to a true expected detection of $\Lambda_{\mathrm{tot}}=5\times 10^{6}$ binaries. We assume that the lensing identification method has a false positive probability of $\alpha=10^{-9}$. We assume different true positive probabilities: $\epsilon=0.4,0.5,0.6,0.8$, corresponding to $k_{0}\equiv\alpha/\epsilon=[2.5,2,1.67,1.25]\times 10^{-9}$. Figure 6 shows the posteriors on the cosmological parameters from an analysis that takes into the effect of contamination. When there is no contamination $k_{0}=0$, we reproduce the results from figure 1. Additionally, data contamination only worsens the precision of our measurement, without causing any systematic biases (see p-p plot in figure 6). Table 2 tabulates the expected precision in the measurement of cosmological parameters with different levels of contamination.
No contamination $k_{0}=2.5$ $k_{0}=2$ $k_{0}=1.67$ $k_{0}=1.25$
@@
Figure 6: Left panel: Posteriors on the cosmological parameters (95% credible regions) when we assume different amounts of data contamination. The legends show $k_{0}\equiv\alpha/\epsilon$ in units of $10^{-9}$, where $\alpha$ ($\epsilon$) is the false (true) positive probability of the lensing identification algorithm. Data contamination only worsens the precision of the posteriors, without causing any systematic biases. Right panel: p-p plot obtained from $\sim 10^{3}$ recovery tests for $\alpha=10^{-9}$ and $\epsilon=0.5$. This indicates no systematic bias in cosmological parameter recovery when we account for contamination.
@@ 7 Conclusions and future work
In this paper we presented a detailed exposition of a statistical method for cosmography from the observation of a large number ($\sim 10^{4}$) of strongly lensed BBH mergers observable by XG detectors. This method, first presented in [33], compares the observed number of strongly lensed GW events and their time delay distribution (between lensed images) with observed events to infer cosmological parameters. We showed that the prospective constraints from the XG detectors are comparable to the current best measurements, but probing a cosmological epoch ($z\sim 1-10$) that is not probed by other observations. This can potentially shed some light on the “tensions” that exist between current cosmological observations [64], some hinting at the failure of the flat $\Lambda$CDM model.
The observed number of lensed events and their time delay distribution will depend on the distribution of GW sources as well as lenses, apart from the cosmological parameters. Properties of the GW source distribution could be measured accurately from the large number ($\sim 10^{6}$) of unlensed GW signals that will dominate the data. Since the distribution of the GW source properties is currently largely unknown, we considered a few different astrophysical models for the same, and showed that the expected constraints don’t vary significantly. We also showed that the statistical uncertainties in the reconstruction of source properties from GW observations will not significantly affect the inference of cosmological parameters, as long as the source redshift distribution can be reconstructed in an unbiased manner. We also showed that the limited ability of our data analysis algorithms to distinguish between lensed and unlensed GW events, resulting in some amount of contamination in the sample of lensed GW events, will not bias our inference. This can be avoided by incorporating the effect of data contamination in our Bayesian likelihood models.
@@
While we focus on BBH mergers in this paper, observations of strongly lensed binaries involving neutron stars could also be powerful. Since BNS could be observed only out to a smaller redshifts ($z\lesssim 2$ even for XG detectors), their lensing optical depth will be smaller compared to BBHs which could be observed out to very large redshifts ($z\sim 10-100$). However, BNS mergers are expected to be large in number ($\sim 10^{6}$ detections per year in XG [65]) and, owing to their long signal duration in the detector band, their properties could be measured with much better accuracy. Furthermore, some of them could also produce observable EM counterparts. Strongly lensed BNSs could be very powerful in probing low-redshift cosmology. We are exploring this in an ongoing work. In conclusion, future observations of strongly lensed GWs will enable several powerful new probes of cosmology.
@@ Acknowledgments:
We thank Otto Akseli Hannuksela for his careful review of the manuscript and useful comments. We are also thankful to Ken K.Y. Ng and Salvatore Vitale for supplying the posterior predicted samples of merger rate density, which were utilised in the analysis conducted in section 4. For useful discussions, we thank Koustav Narayan Maity as well as members of the ICTS Astrophysical Relativity group and the LIGO-Virgo-KAGRA collaboration’s working group on gravitational lensing. Our research is supported by the Department of Atomic Energy, Government of India, under Project No. RTI4001. The numerical calculations were carried out using the Alice cluster at ICTS.